OpenStudy (anonymous):

Given the polynomial below, find F(1). F(x) = x58 - 6x2 + 5

OpenStudy (anonymous):

F(x) = x^58 - 6x^2 + 5

OpenStudy (anonymous):

its 58-1=57

OpenStudy (anonymous):

If you put 1 in for x then you get F(1)=(1)^58-6(1)^2+1. That comes out to 0.